# Code-Aware Prompting: A study of Coverage guided Test Generation in Regression Setting using LLM

### v1

@@ Abstract.
In this paper, we present SymPrompt, a code-aware prompting strategy for LLMs in test generation. SymPrompt’s approach is based on recent work that demonstrates LLMs can solve more complex logical problems when prompted to reason about the problem in a multi-step fashion. We apply this methodology to test generation by deconstructing the testsuite generation process into a multi-stage sequence, each of which is driven by a specific prompt aligned with the execution paths of the method under test, and exposing relevant type and dependency focal context to the model. Our approach enables pretrained LLMs to generate more complete test cases without any additional training. We implement SymPrompt using the TreeSitter parsing framework and evaluate on a benchmark challenging methods from open source Python projects. SymPrompt enhances correct test generations by a factor of 5 and bolsters relative coverage by 26% for CodeGen2. Notably, when applied to GPT-4, symbolic path prompts improve coverage by over 2×2$\times$2 × compared to baseline prompting strategies.
@@
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: The ACM International Conference on the Foundations of Software Engineering; July 15–19, 2024; Porto de Galinha, Brazilbooktitle: FSE ’24: The ACM International Conference on the Foundations of Software Engineering, July 15–19, 2024, Porto de Galinha, Brazilccs: Software and its engineeringccs: Computing methodologies Artificial intelligence
@@ Results.
Refer to caption (a) Focal method. Refer to caption (b) SBST test generations.    Refer to caption (c) LLM test generations.
(d) SymPrompt test generations.
@@
Figure 2. Workflow for generating path constraint prompts. The focal method shown in Figure 0(a) is first parsed and its abstract syntax tree is traversed in preorder. In step \⃝raisebox{-0.9pt}{1}, the traversal first visits the first method statement, normalize_path(path), but does not record any information since it is not a branch constraint. In step \⃝raisebox{-0.9pt}{2}, it then traverses to the first if statement, and records that there is a constraint path.is_dir() that must be satisfied to execute the current path on the AST. It then reaches the return ’directory’ under the first if check, and records that there is an execution path where ’directory’ is returned when path.is_dir() is true. The preorder traversal next visits the if path.is_file(), return ’file’ branch of the AST in \⃝raisebox{-0.9pt}{3} and records a second path where path.is_dir() is false and path.is_file() is true, and the return behavior is ’file’. This traveral continues until in step \⃝raisebox{-0.9pt}{4}, the final return statement is reached, based on an execution path where none of the branch constraints are true. Each collected execution path and return value is then used to construct prompts for test generations that specifies both the path constraints and return behavior for the target test case.
@@ 2. Working Example
SymPrompt. Our approach works in three steps: (i) Collecting Approximate Path Constraints. This step us the core to our approach. Figure 2 shows how path constraints are collected and used to prompt an LLM to generate high coverage tests. Possible execution paths in the focal method are collected by traversing the method’s abstract syntax tree (AST) in preorder and recording branch conditions on each possible execution path, where if a branch is not taken its branch condition is inverted on a given path. When a return statement is reached, the set of branch constraints that need to be satisfied to reach that return statement is recorded, along with the returned value. Each recorded path constraint and return value is then used to generate a prompt, which instructs the LLM to generate a test case targeting the corresponding path.
(ii) Context Construction. Besides the path constraint, each prompt includes includes the focal method signature, which guides the LLM to generate a correct focal method call. We further include additional focal context for generation based on the types and methods that appear in the focal method that exposes relevant data structure and external library dependencies the model may need to reference for effective test generation, as shown in the figure below. For the focal method exists_as, we include the definition of the input parameter _PATH and the external method call normalize_path. One such prompt for our working example is shown in Figure 3.
@@ 3.2.2. Path Minimization
Figure 5. Path minimization algorithmic definition and illustration of how path minimization prevents the number of paths from growing exponentially in the number of branches. A method with $n=3$ if conditions will have $82^{n}=$ 8 possible execution paths, but applying the algorithm in 4(a) reduces the number of paths that need to be tested to at most $n+1=4$, each of which covers a unique branch condition.
@@
Figure 5 illustrates how path minimization reduces the number of collected paths for a simple example method with three sequential if-else branches. The total number of paths in the method is $2\times 2\times 2=8$, but applying Algorithm 4(a) reduces the number of paths to at most 4 (one initial path, $+1$ for each branch). We apply path minimization in Algorithm LABEL:alg:sympath_generation after visitASTNode is called on any if or while statement that introduces new path constraints.
@@
Prior work in test generation with LLMs have demonstrated that including additional context to the focal method definition is beneficial to the quality of test generations (Tufano et al., 2020). For SymPrompt, we focus on exposing two types of context to the model that are beneficial to generating correct test cases: (i) Type context. Context that includes type definitions for focal method parameters guides the model to generate correct test inputs, particularly when the input involves complex data structures. (ii) Dependency Context. We construct the generation focal context selectively to include only type context and dependency context, but also evaluate as a baseline test generation with a local focal context that is constructed greedily by adding lines from the focal file until the model’s input is fully utilized (see Sections 4.1, 4.3). In addition, we construct the test execution context based on the focal context to override any model-generated imports and type definitions. We found that this prevents many common errors, particularly when generating tests with chat-tuned models such as GPT-4 (See Section 4.4).
@@
Figure 7. Example of path constraint prompt construction on a simplified focal method in the tornado project. Each path prompt has three components: \⃝raisebox{-0.9pt}{1} Method Signature. The prompt specifies the test case is for the focal method, including its full signature. \⃝raisebox{-0.9pt}{2} Path Constraints. Next, the prompt specifies what path constraints should be satisfied by the inputs in the given testcase in order to follow the desired execution path. \⃝raisebox{-0.9pt}{3} Return Behavior. The prompt specifies what return behavior, if any, is expected on the specified execution path. This can guide correct generation of assertions for specific return cases.
@@ 4. Evaluation
  • RQ1. Performance Impact. How does SymPrompt effect testing performance over simple test generation prompting methods in a regression setting?
@@ 4.1. RQ1: Performance Improvement
We first evaluate how SymPrompt effects testing performance over simple test generation prompting methods in a regression setting.
@@
Method
Pass@1
FM Call@1
Correct@1
Line Cov.
Branch Cov.
Pass@1
FM Call@1
Correct@1
Line Cov.
Branch Cov.
No-Op Tests
1.00
0.00
0.00
0.33
0.33
1.00
0.00
0.00
0.26
0.36
Baseline Prompt
0.12
0.34
0.03
0.38
0.40
0.12
0.32
0.03
0.32
0.45
SymPrompt
0.41
0.49
0.15
0.48
0.44
0.41
0.35
0.12
0.36
0.35
Pynguin
0.72
0.64
0.68
0.57
SymPrompt Filtered
0.81
1.00
0.81
0.77
0.66
0.83
1.00
0.83
0.71
0.57
@@
Refer to caption (a) Focal method.    Refer to caption (b) SymPrompt test generation.
(c) SBST test generation.
@@ 4.3. RQ3: Design Choices
Table 2. RQ 3. Results (LHS): Ablation results evaluated on 100 randomly sampled focal methods from the benchmark used in RQ1. The ablation results demonstrate that both type and dependency in calling context and path constraint prompts improve correct generations and coverage relative to baseline prompts, but path constraint prompts contribute significantly more to the overall improvement in correct generations and coverage demonstrated by SymPrompt for generations with CodeGen2. RQ 4. Results (RHS): Results with GPT. When evaluating GPT we use as a baseline a two stage describe then generate prompt based on prior work. Although performance is similar for both prompting approaches without calling context, when SymPrompt is used without ablation it gives a relative improvement of 178% in Correct@1 rate and 1.05% relative improvement in coverage over baseline prompts.
@@
Method
Pass@1
FM Call@1
Correct@1
Line Cov.
Branch Cov.
Pass@1
FM Call@1
Correct@1
Line Cov.
Branch Cov.
Baseline Prompt
0.09
0.37
0.04
0.30
0.29
0.14
0.12
0.09
0.36
0.40
Constraints Only
0.27
0.49
0.17
0.49
0.38
0.15
0.15
0.10
0.39
0.43
Context Only
0.16
0.40
0.06
0.42
0.35
0.38
0.28
0.18
0.43
0.47
SymPrompt
0.50
0.65
0.26
0.53
0.42
0.46
0.39
0.25
0.74
0.74

### v2

# Code-Aware Prompting: A Study of Coverage-Guided Test Generation in Regression Setting using LLM

@@ Abstract.
In this paper, we present SymPrompt, a code-aware prompting strategy for LLMs in test generation. SymPrompt’s approach is based on recent work that demonstrates LLMs can solve more complex logical problems when prompted to reason about the problem in a multi-step fashion. We apply this methodology to test generation by deconstructing the testsuite generation process into a multi-stage sequence, each of which is driven by a specific prompt aligned with the execution paths of the method under test, and exposing relevant type and dependency focal context to the model. Our approach enables pretrained LLMs to generate more complete test cases without any additional training. We implement SymPrompt using the TreeSitter parsing framework and evaluate on a benchmark challenging methods from open source Python projects. SymPrompt enhances correct test generations by a factor of 5 and bolsters relative coverage by 26% for CodeGen2. Notably, when applied to GPT-4, SymPrompt improves coverage by over 2×2$\times$2 × compared to baseline prompting strategies.
@@
copyright: rightsretaineddoi: 10.1145/3643769journalyear: 2024submissionid: fse24main-p940-pjournal: PACMSEjournalvolume: 1journalnumber: FSEarticle: 43publicationmonth: 7ccs: Software and its engineeringccs: Computing methodologies Artificial intelligence
@@ Results.
Refer to caption (a) Focal method. Refer to caption (b) SBST test generations.    Refer to caption (c) LLM test generations. Refer to caption (d) SymPrompt test generations.
@@
[…]
@@ 2. Working Example
SymPrompt. Our approach works in three steps: (i) Collecting Approximate Path Constraints. This step is the core to our approach. Figure 2 shows how path constraints are collected and used to prompt an LLM to generate high coverage tests. Possible execution paths in the focal method are collected by traversing the method’s abstract syntax tree (AST) in preorder and recording branch conditions on each possible execution path, where if a branch is not taken its branch condition is inverted on a given path. When a return statement is reached, the set of branch constraints that need to be satisfied to reach that return statement is recorded, along with the returned value. Each recorded path constraint and return value is then used to generate a prompt, which instructs the LLM to generate a test case targeting the corresponding path.
(ii) Context Construction. Besides the path constraint, each prompt includes the focal method signature, which guides the LLM to generate a correct focal method call. We further include additional focal context for generation based on the types and methods that appear in the focal method that exposes relevant data structure and external library dependencies the model may need to reference for effective test generation, as shown in Figure 3. For the focal method exists_as, we include the definition of the input parameter _PATH and the external method call normalize_path. One such prompt for our working example is shown in Figure 3.
@@ 3.2.2. Path Minimization
Figure 5. Path minimization algorithmic definition and illustration of how path minimization prevents the number of paths from growing exponentially in the number of branches. A method with $n=3$ if conditions will have $82^{n}=$ 8 possible execution paths, but applying the algorithm in 5(a) reduces the number of paths that need to be tested to at most $n+1=4$, each of which covers a unique branch condition.
@@
Figure 5 illustrates how path minimization reduces the number of collected paths for a simple example method with three sequential if-else branches. The total number of paths in the method is $2\times 2\times 2=8$, but applying Algorithm 5(a) reduces the number of paths to at most 4 (one initial path, $+1$ for each branch). We apply path minimization after collecting new path constraints from any if or while statement while traversing the focal method AST.
@@
Prior work in test generation with LLMs have demonstrated that including additional context to the focal method definition is beneficial to the quality of test generations (Tufano et al., 2020). For SymPrompt, we focus on exposing two types of context to the model that are beneficial to generating correct test cases: (i) Type context. Context that includes type definitions for focal method parameters guides the model to generate correct test inputs, particularly when the input involves complex data structures. (ii) Dependency Context. Dependency context includes imports and other definitions that are used in the focal method, such as API calls and global data structures. We construct the generation focal context selectively to include only type context and dependency context. In addition, we construct the test execution context based on the focal context to override any model-generated imports and type definitions. We found that this prevents many common errors, particularly when generating tests with chat-tuned models such as GPT-4 (See Section 4.4).
@@
[…]
@@ 4. Evaluation
  • RQ1. Performance Impact. How does SymPrompt affect testing performance over simple test generation prompting methods in a regression setting?
@@ 4.1. RQ1: Performance Improvement
We first evaluate how SymPrompt affects testing performance over simple test generation prompting methods in a regression setting.
@@
Method Pass@1 FM Call@1 Correct@1 Line Cov. Branch Cov. Pass@1 FM Call@1 Correct@1 Line Cov. Branch Cov.
No-Op Tests 1.00 0.00 0.00 0.33 0.33 1.00 0.00 0.00 0.26 0.36
Baseline Prompt 0.12 0.34 0.03 0.38 0.40 0.12 0.32 0.03 0.32 0.45
SymPrompt 0.41 0.49 0.15 0.48 0.44 0.41 0.35 0.12 0.36 0.35
Pynguin - - - 0.72 0.64 - - - 0.68 0.57
SymPrompt Filtered 0.81 1.00 0.81 0.77 0.66 0.83 1.00 0.83 0.71 0.57
@@
Refer to caption (a) Focal method.    Refer to caption (b) SymPrompt test generation. Refer to caption (c) SBST test generation.
@@ 4.3. RQ3: Design Choices
Table 2. RQ 3. Results (LHS): Ablation results evaluated on 100 randomly sampled focal methods from the benchmark used in RQ1. The ablation results demonstrate that both type and dependency in calling context and path constraint prompts improve correct generations and coverage relative to baseline prompts, but path constraint prompts contribute significantly more to the overall improvement in correct generations and coverage demonstrated by SymPrompt for generations with CodeGen2. RQ 4. Results (RHS): Results with GPT. When evaluating GPT we use as a baseline a two stage prompt in which the model is first prompted to describe the method under test, and then generate tests based on both the method definition and the previously generated description. Although performance is similar for both prompting approaches without calling context, when SymPrompt is used without ablation it gives a relative improvement of 178% in Correct@1 rate and 1.05% relative improvement in coverage over baseline prompts.
@@
Method Pass@1 FM Call@1 Correct@1 Line Cov. Branch Cov. Pass@1 FM Call@1 Correct@1 Line Cov. Branch Cov.
Baseline Prompt 0.09 0.37 0.04 0.30 0.29 0.14 0.12 0.09 0.36 0.40
Constraints Only 0.27 0.49 0.17 0.49 0.38 0.15 0.15 0.10 0.39 0.43
Context Only 0.16 0.40 0.06 0.42 0.35 0.38 0.28 0.18 0.43 0.47
SymPrompt 0.50 0.65 0.26 0.53 0.42 0.46 0.39 0.25 0.74 0.74